# Structural response of neutron stars to rapid rotation and its impact on the braking index

### v1

@@ IV.2 Effect of rapid rotation on $n_{\rm EM},n^{Q}_{\rm{GW}}$ and $n^{R}_{\rm{GW}}$
In Sec IV.1, we have shown how the expressions in Eq. 23 reduce to the known values of the braking index in the literature when the rotational effects are ignored by setting all derivatives w.r.t $\Omega$ to zero. However, when the effect of rotation is taken into account the departure of $n_{\rm EM},n^{Q}_{\rm{GW}}$ and $n^{R}_{\rm{GW}}$ from 3, 5 and 7 occurs, which depends on the EOS through the variation of the macroscopic quantities and their derivatives as the function of the $\Omega$.
@@
The response to the rapid rotation depends on the stiffness of the EOS, which is different for different EOSs (refer to Sec. III for detailed discussion), giving rise to different braking index curves. However, the difference in response to the $n_{\rm EM}$ and $n^{R}_{\rm{GW}}$ at frequencies $\lesssim 550$ Hz is negligible for at least these three representative EOS shown in the Fig. 5, unlike in the case of $n^{Q}_{\rm{GW}}$. Both $n_{\rm EM}$ and $n^{R}_{\rm{GW}}$ systematically decrease with the increase in the spin frequency, which indicates at frequencies $\gtrsim 200$ Hz the contribution of $n_{I}$ is significant. The importance of $n_{I}$ can be argued from Fig. 4, where the $R$ is almost constant at $f\sim 200-750$ Hz, indicating $R^{\prime}\sim 0$ which would lead to $n_{\rm EM}\simeq 3$ of Eq. 24 at this frequency range, hence, the deviation seen in $n_{\rm EM}$ from 3 in Fig. 5 can be attributed to $n_{I}$. A similar trend in the variation of $n^{R}_{\rm{GW}}$ can be seen in Fig. 5, which is also attributed to the $n_{I}$. The finite (but small positive) value of $M^{\prime}/M$ and finite (but small negative) value of $\mathcal{J}^{\prime}/\mathcal{J}$ leads to a different slope than the $n_{\rm EM}$ in the initial fall of $n^{R}_{\rm{GW}}$ from the value 7.
@@
The spin frequency evolution of $n^{Q}_{\rm{GW}}$ depends only on $I$ and its derivative which changes appreciably even at $f\gtrsim 200$ Hz, making $n^{Q}_{\rm{GW}}$ always $>5$ (the first term of Eq. 25) and an increasing function of $f$ in the absence of $n_{I}$. Whereas $n_{I}$ is always $<0$, and the nature of its fall depends on the EOS. Therefore, there may exist a range of frequencies for which the $n^{Q}_{\rm{GW}}>5$ in the presence of the $n_{I}$ term as shown in Fig. 5. However, unlike $n_{\rm EM}$ and $n^{R}_{\rm{GW}}$ (at least with the example for these EOS) we find the EOS-dependent variation in $n^{Q}_{\rm{GW}}$ is distinguishable even at the frequencies $\sim 250-300$ Hz. But as the spin frequency increases the variation of all three braking indices for different EOS can be distinguished at the $f\sim 716$ Hz, the spin frequency of PSR J1748--2446ad (maximum spin frequency of a pulsar known till date), an eclipsing binary millisecond pulsar in the globular cluster Terzan 5 (Hessels et al., 2006). However, at much larger frequencies ($f\gtrsim 850$ Hz), we find the large change in radius (see Fig. 5) indicating the size decreases with frequency, which could be an artefact because of forceful fitting of a sphere to a largely deformed star. Hence, the results beyond $f\gtrsim 850$ Hz must be interpreted with care.
@@
Figure 6: This figure captures the dependence of braking index as the function of the rotation frequency on the magnetic field strength $B_{p}$, the ellipticity due to long-lived deformation $\epsilon$ and the r-mode saturation amplitude $\alpha_{\rm S}$. However, we use a constant inclination angle $\alpha=45^{\circ}$. The variation has been studied using three representative EOSs, the red colour lines indicate DDME2 EOS, the blue colour line indicates S271v6 EOS and the green line indicates an EOS from our sample (see Sec. IV.3) for details. The results for different combinations of $B_{p},\epsilon$ and $\alpha_{\rm S}$ have been shown with different line styles. Refer to Sec. IV.3 for detailed discussion and implications.
@@
In the previous Sec. IV.2, we have discussed the rotational effect on all three separate components of the braking index individually in the absence of the other two channels of energy loss. Here we present the result of $\Omega$ dependent braking index taking all three energy loss channels into account for the same set of three representative EOSs used in Sec. IV.2. We further investigate the dependence of the braking index for various physically motivated combinations of parameters of $B_{p},\epsilon$ and $\alpha_{\rm S}$. The results are shown in Fig. 6, where the green colour indicates the results for the EOS obtained from our agnostic EOS model (see Sec. II for details). The same sample EOS has been used in Sec. IV.2 and has been shown with the solid yellow line in Fig. 5. The blue and red colour corresponds to S271v6 and DDME2 EOS, and different line style indicates the different combinations of $B_{p},\epsilon$ and $\alpha_{\rm S}$.
@@
The upper left panel of the Fig. 6 shows the variation of $n(\Omega)$ for neutron stars with a typical value $B_{p}=10^{8}$ G, $\epsilon=10^{-7}$ and for two values of $\alpha_{\rm S}=10^{-1}$ and $10^{-4}$ shown using the solid and dashed line respectively. The justification for using such value of $\epsilon$ and $\alpha_{\rm S}$ has been presented in Sec. IV.1. To arrive at the typical value of $B_{p}$, we find the median value of $B_{p}$ of pulsars with $f>200$ Hz from the ATNF pulsar catalogue Manchester et al. (2005). The overlap of the dashed and solid lines for all the EOSs indicates that the effect of two different parameter sets can not be distinguished. The braking index of $\sim 7$ around $f\sim 200-400$ Hz indicates that the evolution is mostly dominated by the GW emission due to r-mode instability. The MDR loss and GW radiation from the mass quadropolar moment do not have any significant contribution towards the braking index. However, the departure from the value of 7 has a contribution from $n_{I}$ and other frequency-dependent structural quantities (as discussed in Sec. IV.2) which makes $n<7$. Therefore, this figure indicates if the r-mode instability is generated in a typical millisecond pulsar (which is dependent on the spin frequency and the redshifted stellar temperature, for details see Kraav et al. (2024)), with $\alpha_{\rm S}$ ranging between $10^{-4}-10^{-1}$, then the spin-down would be dominated by the GW radiation through the r-mode instability.
At higher frequency $f\sim 716$ Hz the contribution of $n_{I}$ and rotational effects contributing to $n^{R}_{\rm{GW}}$ leads to $76<n<$ 7. At a much higher frequency of $f\gtrsim 850$ Hz, the rapid decline in braking index is due to a sharp fall in the radius as discussed previously in Sec. IV.2 and III.
@@
In the upper right-hand side panel of Fig. 6, we explore the evolution of the braking index in millisecond pulsars in the absence of radiation in GW through r-mode instability. We use $B_{p}=10^{8}$ G, and two separate values of $\epsilon=10^{-7}$ and $10^{-10}$ shown in the solid and dashed line respectively. Both lines overlap with braking index $\sim 3$, the deviation from $n=3$ is due to the rotational effect. This also indicates that the emission loss through GW is insignificant even with a deformation of $\epsilon=10^{-7}$. Our findings are consistent with the current observation limit Abbott et al. (2019b). We find that a large deformation which can generate $\epsilon\gtrsim 10^{-3}$ can result in the emission of GW large enough to impact the rotational evolution of the neutron stars. However, such a high magnitude of ellipticity is unlikely for normal millisecond pulsars Johnson-McDaniel and Owen (2013).
The lower left panel of Fig. 6 shows the evolution of the braking index if a neutron star has an ultra-strong magnetic field $B_{p}=10^{15}$ G, making them fall in the class of magnetars. The main motivation for using such a strong magnetic field is to study the frequency-dependence on the braking index of a newly born millisecond magnetar Dai and Lu (1998), which may form due to the accretion-induced collapse of white dwarfs (Usov, 1992) or a merger of two neutron stars (Rowlinson et al., 2013; Lü and Zhang, 2014). The formation of these objects is believed to manifest as the Gamma Ray bursts (GRBs) Lü et al. (2019). In the lower-left panel of Fig. 6, we find if $\alpha_{\rm S}=10^{-4}$ and $\epsilon=10^{-7}$ at lower frequency $f\sim 200-600$ Hz the braking index is $\sim 3$, indicating the magnetic dipolar radiation dominates the spin-down rate of the star. At higher frequencies, $f\gtrsim 600$ Hz the departure from $n\sim 3$ is due to the rotational effects and is sensitive to the EOS.
@@
When the saturation amplitude of the r-mode is large, that is, $\alpha_{\rm S}=10^{-1}$ (shown in the lower left panel of Fig. 6), we find at low rotation frequency ($f\sim 200$Hz) the braking index is $\sim 3$, but finally rises to $\sim$6 close to $f=716$ Hz. The increase in the braking index from $\sim$ 3.5 to $\sim 6$ is the manifestation of $\Omega^{8}$ dependence on the energy loss via the r-mode. The braking index cannot reach the value of 7 because of the rotational effects where $n_{I}$ is also dominant along with the contribution from $R^{\prime}/R$ and $\mathcal{J}^{\prime}/\mathcal{J}$, where all of them are negative quantities. The behavior of the braking index found here illustrates the importance of incorporating the rotational effects to interpret the spin-down evolution of the remnants produced from the astrophysical transient events discussed above. For example, if the rotational effect is not considered, one can misinterpret a measurement of $n=5$ for a magnetar as the spin-down dominated by the GW emission due to a finite mass quadrupolar moment.
@@
We have further investigated the evolution of the braking index in the parlance of a newborn millisecond magnetar with a much higher value of ellipticity ($\epsilon=10^{-3}$), based on the measurement of Xie et al. (2022). The results are shown in the lower right panel of Fig. 6, we use the same colour convention of EOSs mentioned earlier in the section. The dashed line style indicates the braking index evolution with $B_{p}=10^{15}$, $\epsilon=10^{-3}$ and $\alpha_{\rm S}=10^{-1}$, this is the combination with the highest magnetic field, ellipticity and r-mode saturation amplitude. We find the evolution of the braking index is exactly identical to the parameter set $B_{p}=10^{15}$, $\epsilon=10^{-7}$ and $\alpha_{\rm S}=10^{-1}$ (shown by the solid lines in lower left panel of Fig. 6), which implies that even with 4 orders of magnitude larger ellipticity the evolution is still dominated by the r-mode instability if the saturation amplitude is $10^{-1}$. Whereas, if the $\alpha_{\rm S}\leq 10^{-4}$, we find the evolution of the braking index is dominated by the magnetic dipolar radiation, which is shown by the solid line with the $B_{p}=10^{15}$, $\epsilon=10^{-3}$ and $\alpha_{\rm S}=10^{-4}.$ The dashed-dotted line indicates the evolution of braking index in absence of r-mode but finite deformation of $\epsilon=10^{-3}$ and $B_{p}=10^{15}$, which indicates again the magnetic dipole radiation is the dominant mode of the energy loss from the neutron star.
Figure 7: Evolution of the braking index as the function of rotation frequency. The blue band and the red band show the 90 percentile and 68 percentile of the braking index respectively, which originates from the uncertainties in the EOS computed in Sec. II. The over plot of the yellow and grey band is the braking index of GRB 140903A and GRB 130603B with their $1\sigma$ uncertainty measured from their light curves.
@@ IV.4 EOS uncertainties and the braking index of millisecond magnetar
The long and short GRBs are observed to show an X-ray plateau phase after the prompt emission, which lasts for $\sim 10-1000$ s indicating an ongoing energy injection from the central engine Nousek et al. (2006); O’Brien et al. (2006); Zhang et al. (2006); Rowlinson et al. (2013); Xie et al. (2022); Lü and Zhang (2014). There exists debate on the nature of the central engine but millisecond magnetar is one of the plausible models often used in literature. The spin-down of the magnetar is believed to power the X-ray plateau region of the light curve (Zhang and Mészáros, 2001; Metzger et al., 2011), which is modelled to obtain the braking index (Lasky et al., 2017; Xie et al., 2022). We have shown above that the braking index depends on the structure of the neutron stars which is pronounced at higher frequencies. Given the remnant is a newly born millisecond magnetar with spin period $P\sim 1-3$ ms (Xie et al., 2022) it is expected to have the dependence from EOS on its braking index. Therefore, we compute the range of possible braking index at every given rotational frequency, which gets translated from the uncertainties in the EOS obtained in Sec. II. We have used Eq. 23 to compute the braking index for all the EOS constructed in Sec. II, which has been supplemented with all the frequency-dependent profiles of $M,I,R$ and $\mathcal{J}$ (as discussed in the Sec. III and shown in Fig. 4) and their derivatives. We have used $B_{p}=10^{15}$, $\epsilon=10^{-7}$ and $\alpha_{\rm S}=10^{-4}$, where the values of $\epsilon$ and $\alpha_{\rm S}$ is conservative estimates based on our results of Sec. IV.3. The result of our analysis is shown in Fig. 7. The 90 percentile and the 68 percentile in the braking index as the function of frequency are shown by the blue and red band respectively which have been computed from the weights on the EOS parameter obtained from the Bayesian analysis in Sec. II. The range of the braking index gets wider at higher frequencies due to the contribution of the rotational effects from different EOS of various stiffness. The yellow and the grey horizontal bands are the braking index of GRB 140903A and GRB 130603B with the $1\sigma$ uncertainty measured from their lightcurve by Lasky et al. (2017). The measured braking index passes through our estimated band of braking index indicating if the millisecond magnetar is formed with a typical value of $B_{p}=10^{15}$, $\epsilon=10^{-7}$ and $\alpha_{\rm S}=10^{-4}$ or less, the spin frequency of the central engine can be between $\sim 550-850$ Hz for GRB 140903A and $f\lesssim 650$ Hz for GRB 130603B.

### v2

@@ IV.2 Effect of rapid rotation on $n_{\rm EM},n^{Q}_{\rm{GW}}$ and $n^{R}_{\rm{GW}}$
In Sec IV.1, we have shown how the expressions in Eq. 23 reduce to the known values of the braking index when the rotational effects are ignored by setting all derivatives w.r.t. $\Omega$ to zero. However, when the effect of rotation is taken into account, the departure of $n_{\rm EM},n^{Q}_{\rm{GW}}$ and $n^{R}_{\rm{GW}}$ from 3, 5 and 7 occurs, which depends on the EOS through the variation of the macroscopic quantities and their derivatives as the function of the $\Omega$.
@@
The response to rapid rotation depends on the stiffness of the EOS, which is different for different EOSs (refer to Sec. III for detailed discussion), giving rise to different braking index curves. However, the differences in $n_{\rm EM}$ and $n^{R}_{\rm{GW}}$ for different EOS (at least for the three representative ones shown in Fig. 5) are negligible at frequencies $\lesssim 550$ Hz. Both $n_{\rm EM}$ and $n^{R}_{\rm{GW}}$ systematically decrease with increasing spin frequency, which indicates that at frequencies $\gtrsim 200$ Hz the contribution of $n_{I}$ is significant. The importance of $n_{I}$ can be argued from Fig. 4, where $R$ is almost constant at $f\sim 200-750$ Hz, indicating $R^{\prime}\sim 0$, which would lead to $n_{\rm EM}\simeq 3$ of Eq. 24 in this frequency range. Hence, the deviation seen in $n_{\rm EM}$ from 3 in Fig. 5 can be attributed to $n_{I}$, which becomes increasingly negative with frequency (see bottom right panel of Fig. 5). A similar trend in the variation of $n^{R}_{\rm{GW}}$ can be seen in Fig. 5, which is also attributed to $n_{I}$. The finite but small positive value of $M^{\prime}/M$ and the finite but small negative value of $\mathcal{J}^{\prime}/\mathcal{J}$ lead to a slope different from $n_{\rm EM}$ in the initial fall of $n^{R}_{\rm{GW}}$ from the value 7.
@@
The dependence of $n^{Q}_{\rm{GW}}$ on spin frequency comes only from $I$ and its derivative, which changes appreciably even at $f\gtrsim 200$ Hz, making $n^{Q}_{\rm{GW}}$ always $>5$ (the first term of Eq. 25) and an increasing function of $f$ in the absence of $n_{I}$. Whereas $n_{I}$ is always $<0$, and the nature of its fall depends on the EOS. Therefore, there may exist a range of frequencies for which $n^{Q}_{\rm{GW}}>5$ in the presence of the $n_{I}$ term, as shown in Fig. 5. However, unlike $n_{\rm EM}$ and $n^{R}_{\rm{GW}}$ (at least for these EOS), we find distinguishable EOS-dependent variation in $n^{Q}_{\rm{GW}}$ even at frequencies $\sim 250-300$ Hz. But as the spin frequency increases, the variation of all three braking indices for different EOS can be identified clearly at the $f\sim 716$ Hz, the spin frequency of PSR J1748--2446ad (maximum spin frequency of a pulsar known till date), an eclipsing binary millisecond pulsar in the globular cluster Terzan 5 (Hessels et al., 2006). However, at much larger frequencies ($f\gtrsim 850$ Hz), we find the large change in radius (see Fig. 5), indicating the size decreases with frequency, which could be an artifact because of forceful fitting of a sphere to a largely deformed star. Hence, the results beyond $f\gtrsim 850$ Hz must be interpreted cautiously.
@@
Figure 6: This figure captures the dependence of braking index on the magnetic field strength $B_{p}$, the ellipticity due to long-lived deformation $\epsilon$, and the r-mode saturation amplitude $\alpha_{\rm S}$ as a function of rotation frequency. We use a constant inclination angle $\alpha=45^{\circ}$. The variation has been studied using three representative EOS, the red lines indicate DDME2 EOS, the blue lines indicate S271v6 EOS and the green lines indicate an EOS from our sample (see Sec. IV.3 for details). The results for different combinations of $B_{p},\epsilon$ and $\alpha_{\rm S}$ have been shown with different line styles. Refer to Sec. IV.3 for detailed discussion and implications.
@@
In the previous Sec. IV.2, we have discussed the rotational effect on all three separate components of the braking index individually in the absence of the other two channels of energy loss. There we found that the departure from the expected value of the braking index is purely due to structural change. However, that is not the case when all three energy loss mechanisms are present simultaneously. Rather the variation depends on the strength of the source of energy loss, that is $B_{p},\epsilon,\alpha$ and $\alpha_{\rm S}$. In this section, we present the result of rotational dependence of braking index taking all three energy loss channels into account for the same set of three representative EOSs used in Sec. IV.2. We further investigate the variation in the braking index for various physically motivated combinations of parameters of $B_{p},\epsilon$, and $\alpha_{\rm S}$. The results are shown in Fig. 6, where the green color indicates the results for the EOS obtained from our agnostic EOS model (see Sec. II for details). The same sample EOS has been used in Sec. IV.2 and has been shown with the solid yellow line in Fig. 5. The blue and red colors correspond to S271v6 and DDME2 EOS, and different line style indicates the different combinations of $B_{p},\epsilon$, and $\alpha_{\rm S}$.
@@
The upper left panel of the Fig. 6 shows the variation of $n$ for neutron stars with typical values of $B_{p}=10^{8}$ G, $\epsilon=10^{-7}$ and for two values of $\alpha_{\rm S}=10^{-1}$ and $10^{-4}$ shown using the solid and dashed line respectively. The justification for using such values of $\epsilon$ and $\alpha_{\rm S}$ has been presented in Sec. IV.1. To arrive at the typical value of $B_{p}$, we find the median value of $B_{p}$ of pulsars with $f>200$ Hz from the ATNF pulsar catalogue Manchester et al. (2005). The overlap of the dashed and solid lines for all the EOS implies that the braking index is independent of the value of $\alpha_{\text{S}}$ in this case. The braking index of $\sim 7$ around $f\sim 200-400$ Hz indicates that it is mostly dominated by the GW emission due to r-mode instability. The MDR loss and GW radiation from the mass quadrupole moment do not contribute significantly towards the braking index. However, the departure from the value of 7 is caused by $n_{I}$ and other frequency-dependent structural quantities (as discussed in Sec. IV.2), which makes $n<7$. In summary, this figure suggests that if the r-mode instability is generated in a typical millisecond pulsar (which is dependent on the spin frequency and the redshifted stellar temperature, for details see Kraav et al. (2024)) then the spin-down would be dominated by the GW radiation through the r-mode instability, irrespective of the value of $\alpha_{\text{S}}$. At higher frequency $f\sim 716$ Hz the contribution of $n_{I}$ and rotational effects contributing to $n^{R}_{\rm{GW}}$ leads to $76<n<$ 7. At a much higher frequency of $f\gtrsim 850$ Hz, the rapid decline in braking index is due to a sharp fall in the radius as discussed previously in Sec. IV.2 and III.
@@
In the upper right panel of Fig. 6, we show the braking index as a function of frequency in millisecond pulsars in the absence of GW radiation through r-mode instability. We use $B_{p}=10^{8}$ G, and two separate values of $\epsilon=10^{-7}$ and $10^{-10}$ shown in the solid and dashed line respectively. Both lines overlap with braking index $\sim 3$, the deviation from $n=3$ is due to the rotational effect. It indicates that the emission loss through GW is insignificant even with a deformation of $\epsilon=10^{-7}$. Our findings are consistent with the current observation limit Abbott et al. (2019b). We find that a large deformation which can generate $\epsilon\gtrsim 10^{-3}$ can result in the emission of GW large enough to impact the rotational evolution of the neutron stars. However, such a high magnitude of ellipticity is unlikely for normal millisecond pulsars Johnson-McDaniel and Owen (2013).
@@
The lower left panel of Fig. 6 shows the effect of rotation on the braking index if a neutron star has an ultra-strong magnetic field $B_{p}=10^{15}$ G, making them fall in the class of magnetars. The primary motivation for using such a strong magnetic field is to study the frequency-dependence of the braking index of a newly born millisecond magnetar Dai and Lu (1998), which may form due to the accretion-induced collapse of white dwarfs (Usov, 1992) or a merger of two neutron stars (Rowlinson et al., 2013; Lü and Zhang, 2014). The formation of these objects is believed to manifest as the Gamma Ray bursts (GRBs) Lü et al. (2019). From the figure we observe that if $\alpha_{\rm S}=10^{-4}$ and $\epsilon=10^{-7}$ at lower frequency $f\sim 200-600$ Hz the braking index is $\sim 3$, indicating the magnetic dipolar radiation dominates the spin-down rate of the star. At higher frequencies, $f\gtrsim 600$ Hz, the departure from $n\sim 3$ is due to the rotational effects and is sensitive to the EOS. However, when the saturation amplitude of the r-mode is large, that is, $\alpha_{\rm S}=10^{-1}$, we find at low rotation frequency ($f\sim 200$Hz) the braking index is $\sim 3$, but finally rises to $\sim$6 close to $f=716$ Hz. The increase in the braking index from $\sim 3.5$ to $\sim 6$ is the manifestation of $\Omega^{8}$ dependence on the energy loss via the r-mode. The braking index cannot reach the value of 7 because of the rotational effects where $n_{I}$ is also dominant along with the contribution from $R^{\prime}/R$ and $\mathcal{J}^{\prime}/\mathcal{J}$, where all of them are negative quantities. The behavior of the braking index found here illustrates the importance of incorporating the rotational effects to interpret the spin-down evolution of the remnants produced from the astrophysical transient events discussed above. For example, if the rotational effect is not considered, one can misinterpret a measurement of $n=5$ for a magnetar as the spin-down dominated by the GW emission due to a finite mass quadrupole moment.
@@
We have further investigated the frequency dependence of the braking index in the parlance of a newborn millisecond magnetar with a much higher value of ellipticity ($\epsilon=10^{-3}$), based on the measurement of Xie et al. (2022). The results are shown in the lower right panel of Fig. 6. We use the same colour convention of EOS mentioned earlier in the section. The dashed line style represents the braking index with $B_{p}=10^{15}$, $\epsilon=10^{-3}$, and $\alpha_{\rm S}=10^{-1}$. This is the combination with the highest magnetic field, ellipticity and r-mode saturation amplitude. We find that the braking index depends on the rotation frequency exactly in the same way as the set of values $B_{p}=10^{15}$, $\epsilon=10^{-7}$ and $\alpha_{\rm S}=10^{-1}$ (shown by the solid lines in lower left panel of Fig. 6). It implies that even with 4 orders of magnitude larger ellipticity, the frequency dependence of braking index is still dominated by the r-mode instability if the saturation amplitude is $10^{-1}$. Whereas, if the $\alpha_{\rm S}\leq 10^{-4}$, we find the braking index is dominated by the magnetic dipolar radiation, which is shown by the solid line with the $B_{p}=10^{15}$, $\epsilon=10^{-3}$ and $\alpha_{\rm S}=10^{-4}.$ The dashed-dotted line represents the variation of braking index with frequency in the absence of r-mode but large deformation ($\epsilon=10^{-3}$) and magnetic field ($B_{p}=10^{15}$). The plot again suggests that the magnetic dipole radiation is the dominant mode of energy loss from the neutron star.
[…]
@@ IV.4 EOS uncertainties and the braking index of millisecond magnetar
The long and short GRBs are observed to show an X-ray plateau phase after the prompt emission, which lasts for a few tens to thousands of seconds, indicating an ongoing energy injection from the central engine Nousek et al. (2006); O’Brien et al. (2006); Zhang et al. (2006); Rowlinson et al. (2013); Xie et al. (2022); Lü and Zhang (2014). There exists debate on the nature of the central engine, but a millisecond magnetar is one of the plausible models often used in the literature. The spin-down of the magnetar is believed to power the X-ray plateau region of the light curve (Zhang and Mészáros, 2001; Metzger et al., 2011), which is modeled to obtain the braking index (Lasky et al., 2017; Xie et al., 2022). We have shown above that the braking index of a neutron star is affected by its structural changes due to rotation, especially at higher frequencies. Given the remnant is a newly born millisecond magnetar with spin period $P\sim 1-3$ ms (Xie et al., 2022), its braking index is expected to be different from that of a slowly rotating neutron star. However, the response of a neutron star to rapid rotation also depends on the underlying EOS. To explore the uncertainty in the braking index due to the EOS, we next calculate it at different frequencies for all the EOS constructed in Sec. II. We have used Eq. 23, supplemented by all the frequency-dependent profiles of $M,I,R$ and $\mathcal{J}$ (as discussed in Sec. III and shown in Fig. 4) and their derivatives. Here we set $B_{p}=10^{15}$, $\epsilon=10^{-7}$, and $\alpha_{\rm S}=10^{-4}$, where the values of $\epsilon$, and $\alpha_{\rm S}$ are conservative estimates based on our results of Sec. IV.3. The result is shown in Fig. 7 for a NS of $2M_{\odot}$ baryon mass. The 90% C.I. and the 68% C.I. in the braking index as a function of frequency are shown by the blue and red bands, respectively, which have been computed from the weights on the EOS parameters obtained from the Bayesian analysis in Sec. II. The figure shows that the range of the braking index gets wider as the frequency increases. It happens because the structural changes of a neutron star due to rotation strongly depend on the EOS.
While the band of uncertainty obtained in the braking index captures the uncertainty due to the EOS, uncertainty may arise due to the choice of the baryon mass of the star as well. Therefore, we calculate the braking indices for stars with baryon masses 1.6, 2.0, 2.4, and 2.8 $M_{\odot}$ using the same sample EOS used above in the Sec .IV.2. The results are shown in Fig. 7 by curves of different colors and line styles as described in the caption. We find that the braking index for a massive star falls less rapidly with frequency than a lower-mass star. As a result, the heavier star has a larger braking index at a given frequency than a lighter one. A similar mass dependence is reflected in $n_{I}$. This behavior implies if a massive millisecond magnetar is born from some astrophysical events, it will spin down faster in comparison to a millisecond magnetar born lighter. It is also interesting to note that the uncertainty in the braking index from baryon mass completely masks the uncertainty from the EOS in Fig. 7. Therefore, knowing the EOS and rotation frequency is not enough to infer the braking index of a pulsar. We also need to know its gravitational mass. Please note that one can obtain the baryon mass from the gravitational mass by assuming an EOS if the rotation frequency is known. In other words, knowledge of both the gravitational mass and the braking index of a pulsar is necessary to constrain the EOS. Similar conclusions are expected for other combinations of $B_{p}$, $\epsilon$, and $\alpha_{\rm S}$.
Immediately after the birth of a millisecond magnetar, the object undergoes rapid spin-down, resulting in a significant and time-dependent variation in the braking index. Understanding how the progenitor baryon mass, uncertainties in the EOS, and spin-dependent spin-down evolution influence the energy injection mechanism and the resulting GRB light curve morphology requires a comprehensive and self-consistent numerical treatment. Such an analysis is beyond the scope of the present work.